1) In the market for cotton, the quantity demanded and quantity supplied are expressed as QD = 500 − 25p and QS = 30p − 75 where P is the price per pound of cotton. What is the equilibrium price and equilibrium quantity? Please graph the demand and supply curves, and include the equilibrium price and quantity.
Answer #1

Given the quantity demanded and quantity supplied equations for cotton as , QD = 500 - 25p-----(1) and

QS = 30p - 75-------(2)

Inorder to find the equilibrium price and equilibrium quantity we equate both the equations (1) and (2) as,

500 -25p = 30p - 75

500 + 75 = 30p + 25p

575 = 55p

P = 575 / 55 = 10.45 ~ 10.5//

Substitute the value of P in either equation (1) or (2) we get,

QS = 30 * 10.5 - 75 = 315 - 75 = 240//

Therefore the equilibrium price is 10.5 and the equilibrium quantity is 240.

We can now graph the price and quantity as :-

.S ep 5 Lt 3 ?.5 2​​​​​​